# Finite solvable groups of square-free Euler characteristic

### v1

@@ Abstract.
A characterization is completed for finite groups which act regularly on maps with square-free Euler characteristic, associated with infinite families of square-free Euler characteristic regular maps presented. This is based on a classification of finite groups of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
Keywords: solvable group, regular map, square-free, Euler characteristic
@@ 1. Introduction
Edge-transitive maps are categorized into fourteen classes according to local structures and local actions in [9, 15]. The problem of constructing and classifying important classes of such maps with specific Euler characteristic has attracted considerable attention; refer to [4] - [7] for the Euler characteristic to be a negative prime or product of two primes. We aim towards to a classification of regular maps with square-free Euler characteristic.
@@
It leads us to studying groups of square-free Euler characteristic. This project extends the characterization of almost-Sylow cyclic groups, namely, each Sylow subgroup of odd order is cyclic, and the Sylow $2$-subgroup either is trivial or contains a cyclic subgroup of index 2; refer to Zassenhaus [18], Suzuki [16] and Wong [17].
@@
As usual, for a group $G$ and a subset $\pi$ of prime divisors of $|G|$, denote by $G_{\pi}$ a Hall $\pi$-subgroup of $G$, which is a Sylow subgroup if $\pi$ contains a single prime $p$, and denoted by $G_{p}$. For a positive integer $n$, denote by ${\sf Z}_{n}$ a cyclic group of order $n$, by ${\rm D}_{2n}$ or ${\rm Q}_{2n}$ a dihedral or generalized quaternion group of order $2n$, and by ${\rm A}_{n}$ or ${\rm S}_{n}$ an alternating or symmetric group of degree $n$, respectively. A finite group $G$ is a semidirect product of $A$ by $B$ if $A\lhd G$ and $B<G$ are such that $G/A\cong B$, denoted by G=A:BG=A{:}B. Our first main result is the following theorem, which characterizes finite solvable groups of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
@@
      $K=G_{\{2,3\}}$, explicitly given in Lemma 5.6 (ii) - (iii);
@@
The next theorem shows that automorphism groups of regular maps with square-free Euler characteristic are more restricted.
@@
Let $\mathcal{M}$} be a regular map with square-free Euler characteristic, and let $G={\sf Aut}{\mathcal{M}}$ be solvable. Then G=(A:B):KG=(A{:}B){:}K, where $A$ is abelian, $B$ is nilpotent, $A\cap{\bf Z}(A{:}B)=1$, and $K=G_{2}$ or $G_{\{2,3\}}$ is one of the following groups:
  1. (i)
    $K={\sf Z}_{2}$, ${\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$ or ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$, where $\ell\geqslant 1;$
  2. (ii)
    $K=X$ or $X\times{\sf Z}_{2}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$
  3. (iii)
    $K={\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$.
The paper is organized as follows. After the Introduction, we collect some preliminaries in Section 2, and construct several infinite families of regular maps with square-free Euler characteristic in Section 3. Then, we prove our two main results in Section 4 - 5 and Section 6, respectively.
@@ 2. Preliminaries
Let $\it\Gamma$} be a graph and $\lambda$ be a positive integer. Denote by ${\it\Gamma}^{(\lambda)}$ the graph with each edge of $\it\Gamma$} replaced by $\lambda$ edges.
It is known (refer to [12]) that an arc-transitive (multi-) graph $\it\Gamma$} can be expressed as a coset graph :
${\it\Gamma}={\sf Cos}(G,H,J),\ \mbox{or}\ {\sf Cos}(G,H,HgH)\ \mbox{(for simple graphs)},$
which has vertex set $[G:H]$ and edge set $[G:J]$, such that
    $Hg_{1}$ is incident with $Jg_{0}$ if and only if $g_{1}g_{0}^{-1}\in HJ$; or equivalently,
    $Hg_{1}$, $Hg_{2}$ are adjacent if and only if $g_{2}g_{1}^{-1}\in HgH$ in the simple graph case.
A map $\mathcal{M}$} is a $2$-cell embedding of a graph into a closed surface. Let ${\mathcal{M}}=(V,E,F)$ be a map with vertex set $V$, edge set $E$ and face set $F$, and let $G={\sf Aut}{\mathcal{M}}$. The Euler characteristic of $\mathcal{M}$} is given by Euler formula:
@@
A flag of $\mathcal{M}$} is an incidence triple $(\alpha,e,f)$ with $\alpha\in V$, $e\in E$ and $f\in F$. Given a flag $(\alpha,e,f)$, it is well-known that each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\rm D}_{4}$.
Further, if $\mathcal{M}$} is a flag-regular map (also called a regular map), i.e., $G$ acts regularly on flags, then there exists a triple $(x,y,z)$ of involutions of $G$ which determines the map $\mathcal{M}$} as follows:
$G_{\alpha}=\langle x,y\rangle$, $G_{e}=\langle x,z\rangle\cong{\rm D}_{4}$, $G_{f}=\langle y,z\rangle$.
Meanwhile, the Euler characteristic has the form
$\mbox{$\chi({\mathcal{M}})={|G|\over|G_{\alpha}|}-{|G|\over 4}+{|G|\over|G_{f}|}$}.$
Conversely, if a group $G$ has a triple of involutions $(x,y,z)$ such that
$G=\langle x,y,z\rangle$ and $xz=zx$ $(x\neq z)$,
then there exists a unique regular map ${\mathcal{M}}=(V,E,F)$ with
$V=[G:\langle x,y\rangle],\ E=[G:\langle x,z\rangle],\ F=[G:\langle y,z\rangle],$
and the incidence relation defined by non-empty intersection, such that $G={\sf Aut}{\mathcal{M}}$ acts regularly on flags. Hence, such a triple of involutions is called a regular triple for $G$, and the induced regular map is denoted by ${\mathcal{M}}(G,x,y,z)$. Moreover, the underlying graph of ${\mathcal{M}}(G,x,y,z)$ can be expressed as a coset graph:
${\sf Cos}(G,\langle x,y\rangle,\langle x,z\rangle)$, or ${\sf Cos}(G,H,HzH)$ where $H=\langle x,y\rangle$ (for simple graphs).
In conclusion, each of the following two things determines the other:
    the regular maps arisen from a group $G$;
    the regular triples for $G$.
The following lemma established in [11] characterizes Sylow subgroups of $G={\sf Aut}({\mathcal{M}})$ in the case where $\mathcal{M}$} is with square-free characteristic. We note that if $\chi({\mathcal{M}})$ is square-free, then of course so is $\gcd(\chi({\mathcal{M}}),|{\sf Aut}{\mathcal{M}}|)$.
@@
([11, Lem. 2.2]) Let $\mathcal{M}$} be a map, and let $G={\sf Aut}{\mathcal{M}}$. Assume that $\gcd(\chi({\mathcal{M}}),|G|)$ is square-free. Then each Sylow $p$-subgroup of $G$ has a cyclic or dihedral subgroup of index $p$.
@@ 3. Examples of square-free Euler characteristic regular maps
As Lemma 3.2 shows, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, let $n=4x+2$. Then $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$. According to [8], there are infinitely many positive integers $x$ for which f(x)=(2x+1)(4x1)f(x)=(2x+1)(4x-1) is square-free, so there are infinitely many even values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. $\Box$
@@ 4. $p$-Groups
([10], Cor. 4.9) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. If $G$ has a regular triple, then it is one of the following groups:
  • (i)
    $G={\rm D}_{2^{\ell+1}}$ or ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, $\ell\geqslant 1;$
  • (ii)
    $G={\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, where $\langle a\rangle{:}\langle b\rangle={\rm D}_{2^{\ell+3}}$, $(a,b)^{c}=(a^{2^{\ell+1}+1},b)$, $\ell\geqslant 1$.
@@
According to [13, 5.3.4, pp.136], $G$ is one of the groups stated in this lemma. We only need to verify the statements regarding the automorphism groups of $G$.
(1) The automorphism groups of cyclic groups are well-known; refer to [14, Thm. 7.3].
(2) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ with $\ell>1$. Then by [1, Thm. 3.2], ${\sf Aut}(G)$ is of order
@@
Then ${\sf Aut}(G)$ is solvable, since both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable. Note that the following subgroup
@@
(3) Suppose $G={\sf Z}_{p}\times{\sf Z}_{p}$. Then ${\sf Aut}(G)={\rm GL}(2,p)$, of order $p(p+1)(p-1)^{2}$ with prime divisors less than or equal to $p$. Further, let $M\leqslant{\rm GL}(2,p)$ be a subgroup of order coprime to $2p$, and let $N=M/{\bf Z}(M)$. Then
@@
(4) Suppose that $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$ is non-abelian, $\ell\geqslant 2$. The automorphism group ${\sf Aut}(G)$ is described in [2, Thm. 3.1 & 3.2], by which we have: $|{\sf Aut}(G)|=(p-1)p^{\ell+1}$, ${\sf Aut}(G)$ is solvable with the Sylow $p$-subgroup normal, and $X=\langle\rho\rangle={\sf Z}_{p-1}$ is a Hall $p^{\prime}$-subgroup of $G$, where $\rho$ stabilizes $\langle a\rangle$ and centralizes $\langle b\rangle$, that is, $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$. $\Box$
@@
It is proved in a similar way as the previous lemma. Note that by Lemma 4.3 any prime divisor of $|{\sf Aut}(G_{p})|$ divides p(p1)(p+1)p(p-1)(p+1). $\Box$
@@
By Lemma 4.3 (i), the order of $x$ divides p1p-1. Let $H=\langle a\rangle$ and suppose
@@
Case 2. Let $A_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. Then $r$ does not divide $|B|$. By Lemma 4.3 (ii), it implies that either
@@
Suppose $G_{p}={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}=\langle a\rangle{:}\langle b\rangle$. By Lemma 4.3 (iii), $B$ acts on $\langle a\rangle$ non-trivially and acts on $\langle b\rangle$ trivially. Rewrite $G$ as
@@
Suppose $G_{p}={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$, where $\ell\geqslant 1$. By Lemma 4.3 (ii), it implies that either (i) $G_{p}={\sf Z}_{p}^{2}$ and $B$ is irreducible on $G_{p}$; or (ii) there exists $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$. For the former, G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$. For the former, if $B$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $C=G_{p}\times A$, $D=B$. Otherwise, $B$ acting non-trivially on $G_{p}$ implies $B$ acting non-trivially on some one of $\langle a\rangle$ and $\langle b\rangle$, say $\langle a\rangle$. Then similar to the last paragraph, let $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$. $\Box$
@@
  • (ii)
    $G={\sf Z}_{2}^{2}{:}H$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H)$, ${\rm Q}_{8}{:}H$, or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}H)$, where $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, ${\sf Z}_{2}^{2}{:}H\cong{\rm A}_{4}$ or ${\rm S}_{4}$, and ${\rm Q}_{8}{:}H\cong{\rm SL}(2,3)$ or ${\rm GL}(2,3)$.
@@
(1) ${\bf O}_{2}(\overline{K})\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$; or (2) ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{3}$, $H={\sf Z}_{7}$ or ${\sf Z}_{3}$.
Further, if (2) holds, $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{p}$ with $p=3$ or $7$, then ${\sf Z}_{2}^{3}$ is a minimal normal subgroup of $\overline{K}$, and so $\overline{F_{2}}=F_{2}$ as a normal subgroup of $\overline{K}$ must be identity. Consequently, we have $F=F_{p}$, ${\sf Z}_{2}^{3}{:}{\sf Z}_{p}=K/F_{p}\lesssim{\sf Out}(F_{p})$, which does not hold for the $p$-groups appearing in Lemma 4.3.
Therefore, we have that (1) holds, and since $\overline{K}={\bf O}_{2}(\overline{K}){:}H$ is known, $F_{2}=\overline{F_{2}}$ as a normal 2-subgroup of $\overline{K}$ is also known. We then analyze all the possibilities.
@@
Table 1. Case: $F_{2}={\bf O}_{2}(\overline{K})$
@@
(1.3) ${\sf Z}_{3^{\ell}}\times{\rm S}_{4},$ ${\sf Z}_{2.3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})$ ${\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})$
(1.4) ${\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$, ${\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{3}\times({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{12}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$
@@
Table 2. Case: $F_{2}<{\bf O}_{2}(\overline{K})$
@@
We are now ready to prove the first main theorem of this paper.
@@
In this section we focus on the automorphism groups of regular maps.
The following lemma shows that the class of groups with regular triples is closed under taking quotients.
@@
Let $G$ be a finite group which has a regular triple, and let $\overline{G}$ be a non-identity factor group of $G$. Then either $\overline{G}$ has a regular triple, or $\overline{G}$ is ${\sf Z}_{2}$ or a dihedral group. In particular, $\overline{G}$ is of even order.

### v2

# Finite groups and arc-transitive maps of square-free Euler characteristic

@@ Abstract.
A characterization is completed for finite groups acting arc-transitively on maps with square-free Euler characteristic, associated with infinite families of regular maps of square-free Euler characteristic presented. This is based on a classification of finite groups of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
Keywords: solvable group, arc-transitive map, regular map, Euler characteristic
@@ 1. Introduction
Edge-transitive maps are categorized into fourteen classes according to local structures and local actions in [9, 16], among which five are arc-transitive. The problem of constructing and classifying important classes of such maps with specific Euler characteristic has attracted considerable attention; refer to [4] - [7] for the Euler characteristic to be a negative prime or product of two primes. We aim towards to a classification of arc-transitive maps with square-free Euler characteristic.
@@
It leads us to studying groups of square-free Euler characteristic, which in fact extends the characterization of almost-Sylow cyclic groups, namely, each Sylow subgroup of odd order is cyclic, and the Sylow $2$-subgroup either is trivial or contains a cyclic subgroup of index 2; refer to Zassenhaus [19], Suzuki [17] and Wong [18].
@@
As usual, for a group $G$ and a subset $\pi$ of prime divisors of $|G|$, denote by $G_{\pi}$ a Hall $\pi$-subgroup of $G$, which is a Sylow subgroup if $\pi$ contains a single prime $p$, and denoted by $G_{p}$. For a positive integer $n$, denote by ${\sf Z}_{n}$ a cyclic group of order $n$, by ${\rm D}_{2n}$ or ${\rm Q}_{2n}$ a dihedral or generalized quaternion group of order $2n$, and by ${\rm A}_{n}$ or ${\rm S}_{n}$ an alternating or symmetric group of degree $n$, respectively. A finite group $G$ is a semidirect product of $A$ by $B$ if $A\lhd G$ and $B<G$ are such that $G/A\cong B$, denoted by G=A:BG=A{:}B. Our first main result is the following theorem.
@@
      $K=G_{\{2,3\}}$, explicitly given in Lemma 5.6 (2) - (3);
@@
The next theorem shows that arc-transitive automorphism groups of maps with square-free Euler characteristic are more restricted.
@@
Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be solvable and arc-transitive on $\mathcal{M}$}. Then G=(A:B):KG=(A{:}B){:}K such that $A$ is abelian with $A\cap{\bf Z}(A{:}B)=1$, $B$ is nilpotent, and either $K=G_{2}$ is ${\sf Z}_{2}$ or one of the groups given in Proposition 4.2, or one of the following holds:
  • (1)
    $\mathcal{M}$} is $G$-regular, $K=G_{\{2,3\}}$ is one of $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$
  • (2)
    $\mathcal{M}$} is $G$-vertex-reversing, $K=G_{\{2,3\}}$ is one of the groups in Table 1 and Table 2, corresponding to cases (1.1), (1.2), (1.5), (1.6), (2.4), (2.5);
  • (3)
    $\mathcal{M}$} is $G$-vertex-rotary, $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$, or $K=G_{\{2,3\}}$ is one of ${\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}}$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})$, ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})$, or one of the groups in Table 1 and Table 2, excluding cases (1.5), (1.6).
The proofs of the above two results are given in Section 4 - 5 and Section 6. Before that, we construct several interesting examples of regular maps with square-free Euler characteristic in Section 3, from which the following result follows.
Theorem 1.3.
There are infinite families of regular maps $\mathcal{M}$} with underlying graphs ${\bf C}_{n}^{(n)}$ or ${\bf C}_{n}\square{\bf C}_{n}$, and with Euler characteristic $\chi({\mathcal{M}})=-n(n-3)$ or n(n2)-n(n-2), respectively. In particular, there are infinitely many of them with $\chi({\mathcal{M}})$ square-free.
@@ 2. Preliminaries
Let ${\it\Gamma},\Sigma$ be two graphs and $\lambda$ be a positive integer. Denote by ${\it\Gamma}^{(\lambda)}$ the graph with each edge of $\it\Gamma$} replaced by $\lambda$ edges. Denote by ${\it\Gamma}\square\Sigma$ the Cartesian product of $\it\Gamma$} and $\Sigma$, which has vertex set $V({\it\Gamma})\times V(\Sigma)$, where $(x_{1},y_{1})$ is adjacent to $(x_{2},y_{2})$ if and only if $x_{1}=y_{1}$ and $x_{2}$ is adjacent to $y_{2}$, or $x_{1}$ is adjacent to $y_{1}$ and $x_{2}=y_{2}$.
A map ${\mathcal{M}}=(V,E,F)$ is a $2$-cell embedding of a graph into a closed surface, with vertex set $V$, edge set $E$ and face set $F$. The Euler characteristic of $\mathcal{M}$} is given by Euler formula:
@@
A flag of $\mathcal{M}$} is an incidence triple $(\alpha,e,f)$ with $\alpha\in V$, $e\in E$ and $f\in F$. Let $G\leqslant{\sf Aut}({\mathcal{M}})$, and let $(\alpha,e,f)$ be a flag. It is well-known that each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\rm D}_{4}$.
A map $\mathcal{M}$} is called a $G$-arc-transitive map, if $G$ acts transitively on arcs (directed edges) of $\mathcal{M}$}. Arc-transitive maps are divided into five types in [9, 16], for which recent papers [12, 13] provide a new description: letting $\mathcal{M}$} be a $G$-arc-transitive map, then one of the following holds:
(1) $\mathcal{M}$} is a $G$-regular map (type 1), which is determined a regular triple $(x,y,z)$ for $G$ satisfying:
$\mbox{$G=\langle x,y,z\rangle$, $|x|=|y|=|z|=2$ and $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$}.$
In this case, we have:
$G={\sf Aut}({\mathcal{M}}),\,G_{\alpha}=\langle x,y\rangle,\,G_{e}=\langle x,z\rangle,\,G_{f}=\langle y,z\rangle,$
and the underlying graph of $\mathcal{M}$} can be expressed as a coset graph ${\sf Cos}(G,H,HzH)$, where $H=\langle x,y\rangle$ (refer to [12].)
(2) $\mathcal{M}$} is a $G$-vertex-reversing map (type $2^{\ast}$ or $2^{P}$), which is determined by a reversing triple $(x,y,z)$ for $G$ satisfying: $G=\langle x,y,z\rangle$ and $|x|=|y|=|z|=2$.
(3) $\mathcal{M}$} is a $G$-vertex-rotary map (type $2^{*}{\rm ex}$ or $2^{P}{\rm ex}$), which is determined by a rotary pair $(\alpha,z)$ for $G$ satisfying: $G=\langle\alpha,z\rangle$ and $|z|=2$.
The following lemma established in [11] characterizes Sylow subgroups of ${\sf Aut}({\mathcal{M}})$ in the case where $\mathcal{M}$} is with square-free characteristic. We note that if $\chi({\mathcal{M}})$ is square-free, then of course so is $\gcd(\chi({\mathcal{M}}),|{\sf Aut}{\mathcal{M}}|)$.
@@
([11, Lem. 2.2]) Let $\mathcal{M}$} be a map such that $\gcd(\chi({\mathcal{M}}),|{\sf Aut}({\mathcal{M}})|)$ is square-free. Then each Sylow $p$-subgroup of ${\sf Aut}({\mathcal{M}})$ has a cyclic or dihedral subgroup of index $p$.
@@ 3. Examples of square-free Euler characteristic regular maps
By Lemma 3.2, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, let $n=4x+2$. Then according to [8], there are infinitely many positive integers $x$ for which f(x)=(2x+1)(4x1)f(x)=(2x+1)(4x-1) is square-free, so there are infinitely many even values of $n$ such that $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$ is square-free. $\Box$
@@ 4. $p$-Groups
([10], Thm. 1.1) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. Then:
  • (1)
    If $G$ has a reversing triple, then $G={\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$ or ${\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$, where $\ell\geqslant 1;$ further, if $G$ has a regular triple, then $G\neq{\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4};$
  • (2)
    If $G$ has a rotary pair, then $G={\sf Z}_{2^{\ell+1}}$, ${\sf Z}_{2^{\ell}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+2}}$ or ${\sf Z}_{2^{\ell+2}}{:}{\sf Z}_{2}$, $\ell\geqslant 1$.
@@
According to [14, 5.3.4, pp.136], $G$ is one of the groups stated in this lemma. We only need to verify the statements regarding the automorphism groups of $G$.
(i) The automorphism groups of cyclic groups are well-known; refer to [15, Thm. 7.3].
(ii) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ with $\ell>1$. Then by [1, Thm. 3.2], ${\sf Aut}(G)$ is of order
@@
Now both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable, so is ${\sf Aut}(G)$. Note that the following subgroup
@@
(iii) Suppose $G={\sf Z}_{p}\times{\sf Z}_{p}$. Then ${\sf Aut}(G)={\rm GL}(2,p)$ is of order $p(p+1)(p-1)^{2}$, with prime divisors less than or equal to $p$. Further, let $M\leqslant{\rm GL}(2,p)$ be a subgroup of order coprime to $2p$, and let $N=M/{\bf Z}(M)$. Then
@@
(iv) Suppose that $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$ is non-abelian, $\ell\geqslant 2$. The automorphism group ${\sf Aut}(G)$ is described in [2, Thm. 3.1 & 3.2], by which we have: $|{\sf Aut}(G)|=(p-1)p^{\ell+1}$, ${\sf Aut}(G)$ is solvable with the Sylow $p$-subgroup normal, and $X=\langle\rho\rangle={\sf Z}_{p-1}$ is a Hall $p^{\prime}$-subgroup of $G$, where $\rho$ stabilizes $\langle a\rangle$ and centralizes $\langle b\rangle$, that is, $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$. $\Box$
@@
By Lemma 5.1, $G=G_{p}{:}G_{q}$. Since $G$ is solvable, we obtain that $G_{q}/{\bf C}_{G_{q}}(G_{p})\cong G/(G_{p}{\bf C}_{G_{q}}(G_{p}))=G/(G_{p}{\bf C}_{G}(G_{p}))\leq{\sf Out}(G_{p})$. By Lemma 4.3, any prime divisor of $|{\sf Out}(G_{p})|$ divides $p(p^{2}-1)$. However, $q\nmid p(p^{2}-1)$ by our assumption, and we conclude that $G_{q}/{\bf C}_{G_{q}}(G_{p})=1$, so that $G_{q}={\bf C}_{G_{q}}(G_{p})$ and $G=G_{p}\times G_{q}$. $\Box$
@@
By Lemma 4.3 (1), the order of $x$ divides p1p-1. Let $H=\langle a\rangle$ and suppose
@@
Case 2. Let $A_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. Then $r$ does not divide $|B|$. By Lemma 4.3 (2), it implies that either
@@
Suppose $G_{p}={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}=\langle a\rangle{:}\langle b\rangle$. By Lemma 4.3 (3), $B$ acts on $\langle a\rangle$ non-trivially and acts on $\langle b\rangle$ trivially. Rewrite $G$ as
@@
Suppose $G_{p}={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$, where $\ell\geqslant 1$. By Lemma 4.3 (2), it implies that either (i) $G_{p}={\sf Z}_{p}^{2}$ and $B$ is irreducible on $G_{p}$; or (ii) there exists $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$. For the former, G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$. For the latter, if $B$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $C=G_{p}\times A$, $D=B$. Otherwise, $B$ acting non-trivially on $G_{p}$ implies $B$ acting non-trivially on some one of $\langle a\rangle$ and $\langle b\rangle$, say $\langle a\rangle$. Then similar to the last paragraph, let $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$. $\Box$
@@
  • (2)
    $G={\sf Z}_{2}^{2}{:}H$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H)$, ${\rm Q}_{8}{:}H$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}H)$, where $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, ${\sf Z}_{2}^{2}{:}H\cong{\rm A}_{4}$ or ${\rm S}_{4}$, and ${\rm Q}_{8}{:}H\cong{\rm SL}(2,3)$ or ${\rm GL}(2,3)$.
@@
(i) ${\bf O}_{2}(\overline{K})\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$; or (ii) ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{3}$, $H={\sf Z}_{7}$ or ${\sf Z}_{3}$.
Further, if (ii) holds, $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{p}$ with $p=3$ or $7$, then ${\sf Z}_{2}^{3}$ is a minimal normal subgroup of $\overline{K}$, and so $\overline{F_{2}}=F_{2}$ as a normal subgroup of $\overline{K}$ must be identity. Consequently, we have $F=F_{p}$, ${\sf Z}_{2}^{3}{:}{\sf Z}_{p}=K/F_{p}\lesssim{\sf Out}(F_{p})$, which does not hold for the $p$-groups appearing in Lemma 4.3.
Therefore, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$ is given in (i). We then analyze all the possibilities for $F_{2}=\overline{F_{2}}$, which a normal 2-subgroup of $\overline{K}$.
@@
Table 1. Case 1: $F_{2}={\bf O}_{2}(\overline{K})$
@@
(1.3) ${\sf Z}_{3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{2.3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})$ ${\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})$
(1.4) ${\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{3}\times({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{12}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$
@@
Table 2. Case 2: $F_{2}<{\bf O}_{2}(\overline{K})$
@@
Now we are ready to prove Theorem 1.1.
@@
In this section we focus on the arc-transitive automorphism groups of maps. The following lemma shows that the class of such groups is closed under taking quotients.
@@
Let $G$ be a finite group, with $\overline{G}$ a non-identity factor group. Then:
  • (1)
    If $G$ has a regular triple (resp. reversing triple), then either $\overline{G}$ has a regular triple (resp. reversing triple), or $\overline{G}$ is ${\sf Z}_{2}$ or a dihedral group. In particular, $\overline{G}$ is of even order.
  • (2)
    If $G$ has a rotary pair, then either $\overline{G}$ has a rotary pair, or $\overline{G}$ is cyclic. In particular, if $\overline{G}$ is of odd order, then it is cyclic.